# A Physics Preserving Neural Network Based Approach for Constitutive Modeling of Isotropic Fibrous Materials

### v4

@@ 1 Introduction
Over the past decade, machine learning models have become an increasingly important part of the engineering design process, due to promising accuracy in predicting complex deformation fields and in constitutive modeling of nonlinear materials [17, 18, 19, 20, 21, 22]. Mathematically, a neural network is generally an involved composite function that contains controlled nonlinearity and a large number of free parameters that are calibrated to model a plethora of phenomena. The mathematical formulation is generally discussed in terms of ’neural network architecture’ and can be chosen so that the underlying functional form is convex (e.g., Input Convex Neural Network (ICNN) [23], which is particularly conducive to continuum mechanics based material model formulation). Once trained, these surrogate models have computational efficiency comparable to traditional hyperelastic models while being more capable of handling nuances of complex microstructural deformations. In this paper, we adopt this framework to model fibrous network material using an extensive dataset obtained using MuMFiM.
@@ 2.1.1 Macroscale Problem
$\delta P=\delta P^{\text{int}}-\delta P^{\text{ext}}=0\quad\forall\delta\mathbf{v}\in\mathscr{v}$ (1)
@@
Here, $\mathscr{v}$ is the set of kinematically admissible velocity fields, ${\bf\it\sigma}$ is the Cauchy stress, $\mathbf{D}$ is the symmetric rate of deformation tensor, $\mathbf{t}$ is the surface traction, $V$ is the total volume of the body, and $\partial V$ is the surface of the domain.
@@ 2.1.2 Downscaling
where $vb{x}^{\text{s}}$ is the subscale coordinate, $\mathbf{F}^{\text{M}}$ is the macroscopic deformation gradient at the integration point, and $\tilde{\mathbf{x}}^{\text{s}}$ are the higher-order terms that are typically described as fluctuations. To proceed, we make the following choices, which can be interpreted as orthogonality of the mean and fluctuating fields [24, 25].
@@ 2.1.4 Subscale
In this work, we use Delaunay fiber networks. The network construction procedure proceeds as follows: First, a generation box is seeded with points using a uniform random distribution. A Delaunay triangularization is then constructed from the initial seed points. To avoid boundary effects, the generation box is trimmed so that the new edge length is half of the one used for the generation box. New nodes are inserted at the intersection of the trimming box and any crossing fibers. The fiber segments that remain outside the trimming box are deleted. This process will yield fiber networks with an average connectivity of $z\approx 14$. This is well above the isostaticity limit for 3D truss networks. The constitutive model for each fiber is given by:
@@
Figure 2: Schematic of the input convex neural network we implement. The input vector consists of invariants of the right Cauchy-Green deformation tensor ($\mathbf{C}$), and the output is the predicted strain energy density ($\mathcal{W}$). The components of true stress ($\mathbf{\sigma}$), and stiffness tensor ($\mathop{\mathbb{D}}$), are computed using automatic differentiation of the trained model.
@@ 2.2 Machine Learning Model
The hyperelastic formulation entails finding a non-negative, non-decreasing, and convex scalar energy density function ($\mathcal{W}$), in terms of deformation-related quantities [32, 33]. The stress and stiffness tensor components are then found through appropriate first and second-directional derivatives of $\mathcal{W}$. The convexity condition implies the machine learning algorithm must be chosen or designed appropriately [23, 34]. Here, we employ the Input Convex Neural Network (ICNN) model, which guarantees that the overall mathematical form of the neural network will be convex for the proper choice of input and output quantities. The ICNN network was originally pioneered in [23]. Various authors have used the ICNN as a basis for hyperelastic material modeling [34, 17, 33, 35, 36]. A schematic of the neural network model we implemented is shown in Figure 2.
@@
ICNN is a particular form of fully connected neural network with restrictions on activation functions and weights. The convexity of the model is ensured by strictly adhering to (i) convex activation functions, and (ii) non-negative weights. While the first requirement is trivial, non-negative weights limit the scope of the neural network when modeling complex phenomena [23]. To counter, a number of pass-through layers without restrictions on weights or biases are introduced that connect the input to the hidden layer neurons. These pass-through layers amount to affine mapping of input to hidden layer neurons. Therefore, if the ICNN output is a convex function of the input, the convexity condition on the strain energy density function is satisfied by construction.
Here, we aim to map the principal invariants of the Right Cauchy-Green deformation tensor to the deformed material’s strain energy density through ICNN. Following [36], the input layer of ICNN consists of modified invariants of the right Cauchy-Green deformation tensor ($I_{1}-3$, $I_{2}-3$, and $I_{3}-1$). This choice helps to normalize the input values. We employ the Soft Plus activation function defined as $g(x)=1/\beta\log(1+\beta\exp(x))$ with $\beta=1$, a choice inspired by [17, 34]. The function is non-negative, non-decreasing, and convex in $x$. A second function ($f$), of similar characteristics, is applied over the weights to fulfill the requirements for positive weights. Here, we choose $f=g$ without loss of generality. Therefore, the resulting general form of the strain energy density function is non-negative, non-decreasing, and convex by construction.
@@
This model, however, does not guarantee vanishing strain energy at zero strain (i.e., $\mathcal{W}(\mathbf{F}=\mathbf{I})=0$), due to the nature of activation functions. The optimization algorithm can overcome this shortcoming by adjusting the biases in the pass-through layers. Figure 4 indicates that the requirement is approximately satisfied. An alternative approach can be introducing standalone bias terms to act as correction factors, for example, in [17].
@@
Figure 3: Mean squared error (loss) of ICNN during training. The average errors associated with strain energy density, stress, and stiffness tensors are shown in the first, second, and third columns, respectively. The top row shows the loss for the training dataset, and the bottom row represents the corresponding results for the test dataset. The legend $\mathcal{L}^{k}:=\sum_{i=0}^{i=k}H_{i}$ indicates the variant of the loss function used to train ICNN, with $k$ being the highest order of derivative included in the composite loss function.
@@
The ICNN we train contains five fully connected hidden layers, each having four neurons in the primary branch. The pass-through layers connect the input vector to these hidden layer neurons, giving a total of 164 trainable parameters. Mathematically, the output of the neurons in the $i$-th hidden layer is $\mathbf{z}^{(i)}=g(f(\mathbf{A}^{(i)})\mathbf{z}^{(i-1)}+\mathbf{A}^{(i)}_{p}\mathbf{x}+\mathbf{b}^{(i)}_{p})$. Here, $\mathbf{x}$ is the input vector to the model, and $A^{(i)}$ is the weight matrix for the $i$-th layer of the primary branch. The final two terms ($A^{(i)}_{p}$ and $b^{(i)}_{p}$) represent the weight matrix and bias vector due to the pass-through connection to the $i$-th layer, respectively. The model is trained on 91,590 samples, with an 80/20 split for the train/test dataset. We use Adams solver for parameter optimization without any weight decay [37]. The model parameters are initialized using Kaiming uniform distribution [38]. We use the mean squared error $\text{MSE}(y)=(y_{\text{pred}}-y_{\text{true}})^{2}$ as the template for the loss function.
@@
In the original and most subsequent implementations of ICNN, the loss function is based on the output variable only. For our purpose, this translates to hyperparameter optimization based on strain energy density. However, due to the large number of trainable parameters, the prediction accuracy for energy does not generally translate to its derivative quantities such as stress and stiffness, which are necessary for FE simulations. To increase the prediction accuracy of the derivatives, we implement the so-called Sobolev training protocol such that the model is trained on the target output and its derivatives concurrently [39]. Here, we include both the first and second derivatives in the ICNN training procedure as follows. For the $i$-th batch iteration within an epoch, in addition to the loss in energy ($H_{0}=\text{MSE}(\mathcal{W})$), we compute the component-average loss in stress ($H_{1}=\langle MSE(\mathbf{\sigma})\rangle$) and stiffness ($H_{2}=\langle\text{MSE}(\mathop{\mathbb{D}})\rangle$) through automatic differentiation of the ICNN (updated in the $i-1$-th iteration). The total loss per batch is computed as $\mathcal{L}^{k}=\sum_{j=0}^{j=k}H_{j}$, with $k=2$ being the highest order of derivative considered. The model parameters are then updated based on the total loss, and the cycle is repeated. The learning rate is adjusted automatically at the end of an epoch based on the mean loss such that the learning rate decreases by a factor of ten if the loss exhibits a plateau for five consecutive epochs. We use a moderately large batch size, so the loss (especially in the gradients) approximated from a batch is representative of the population. Finally, we note that the construction of neural network determines the convexity and smoothness of the model. Sobolev training is an add-on intended for accurate prediction of higher-order quantities only.
@@ 2.2.1 Training Data Generation
If a constant number of samples is used within each ellipsoidal shell that is represented by a given base stretch $\mathbf{U}^{*}_{i}$, then each data point will cover a varying volume in phase space. To avoid this problem, the number of rotation samples constructed for each set of base stretches is scaled by the surface area of the ellipsoid, with $\lambda_{1},\lambda_{2},\lambda_{3}$ forming the ellipsoid axes. This will ensure that, in a statistical sense, each sampled point will cover a constant volume in phase space, which is critical to avoid oversampling and associated overfitting. These sampled deformations are used as inputs to our GPU accelerated network solver in MuMFiM to generate the neural network training and test dataset.
@@ 2.2.2 Model Availability
Figure 4: True (MuMFiM) vs ICNN-predicted strain energy density for various deformation gradients ($F$) in test dataset using ICNN trained with $\mathcal{L}^{2}$ loss function. The dashed red line indicates the ideal fit, and scatters indicate predictions. The value of normalized mean squared error is 0.15% for this dataset.
@@ 3 Results and Discussions
In this section, we first present the model’s predictive capability and the improvements brought forth by the Sobolev training protocol. We then present a case study involving mechanical deformation of the Facet Capsule Ligament, simulated using the trained model. Finally, we compare the output to the results of MuMFiMbased large-scale numerical simulation to comment on the efficacy of the current approach.
@@ 3.1 Sobolev Training of ICNN
We first discuss the overall predictive accuracy of the input convex neural network (ICNN) in terms of the average loss of desired quantities. This energy, stress, and stiffness result is shown in the first, second, and third columns of Figure 3. We compare three variants of the model emanating from the choice of the composite loss function (i.e., $k=0,1,2$ in $\mathcal{L}^{k}=\sum_{j=0}^{j=k}H_{j}$), indicating the highest derivatives included in the loss function.
The first column of Figure 3 shows the loss in strain energy density for different choices of the overall loss function. Expectedly, the choice of loss function has a negligible effect on the predictive accuracy of strain energy density. The derivative quantities (such as stress), however, benefit from including $H_{1}$ in the loss function (curves corresponding to $\mathcal{L}^{1}$). This result is shown in the second column of Figure 3. The additional information passed to the optimizer also improved the predictive accuracy of stiffness, as shown in the final column of the figure. Naturally, including the loss in stiffness in the overall loss function (curves corresponding to $\mathcal{L}^{2}$) improved the predictive accuracy of the stiffness for both the training and test datasets. Comparing the loss function variants, it is evident that including higher-order derivatives of the output function improved the accuracy of the derivative quantities at any given epoch. The substantial improvement in stiffness prediction accuracy is especially beneficial for the downstream finite element algorithm, which would be a typical use case of the ICNN-based model developed here. Consequently, subsequent discussions focus on the $\mathcal{L}^{2}$ variant only.

### v5

@@ 1 Introduction
Over the past decade, machine learning models have become an increasingly important part of the engineering design process due to promising accuracy in predicting complex deformation fields and in constitutive modeling of nonlinear materials [17, 18, 19, 20, 21, 22]. Mathematically, a neural network is generally an involved composite function that contains controlled nonlinearity and a large number of free parameters that are calibrated to model a plethora of phenomena. The mathematical formulation is generally discussed in terms of “neural network architecture” and can be chosen so that the underlying functional form is convex (e.g., Input Convex Neural Network (ICNN) [23], which is particularly conducive to continuum mechanics based material model formulation). Once trained, these surrogate models have computational efficiency comparable to traditional hyperelastic models while being more capable of handling nuances of complex microstructural deformations. In this paper, we adopt this framework to model fibrous network material using an extensive dataset obtained using MuMFiM. Our main goal is to accelerate numerical simulations of multiscale models rather than to fit experimentally obtained data as is typical in the literature, for example, [22, 17, 24, 25]. This choice affords a different set of modeling constraints where we seek to obtain the highest quality downstream finite element solution even when it requires data that would be difficult or impossible to obtain experimentally—like full-field stress and material stiffness measurements.
The desire for polyconvex energy functions has been discussed at length in the constitutive modeling literature as it guarantees the ellipticity of the finite element problem and the existence of a minimizer of the elastic free energy and is consistent with objectivity unlike strict convexity [26, 27]. There are two main ways in which this constraint is included in neural networks. First, the so-called “Physics-Informed” approach where convexity is added as a constraint to the loss function. This is the approach that is taken in [24, 22].
An alternative approach is to strictly enforce polyconvexity through the network architecture. Simply put, the network architecture must guarantee convexity of the deformation gradient ($\mathbf{F}$), cofactor of $\mathbf{F}$, and the determinant of $\mathbf{F}$ or accounting for frame indifference convexity with respect to the components of the right Cauchy-Green deformation tensor ($\mathbf{C}$, $\text{cof}\,\mathbf{C}$, and $\det\mathbf{C}$ ) [28]. This approach is taken in [29, 30, 25]. We also use this approach here, preferring strongly enforced constraints to weakly enforced constraints.
@@ 2.1.1 Macroscale Problem
$\delta P=\delta P^{\text{int}}-\delta P^{\text{ext}}=0\quad\forall\delta\mathbf{v}\in{\bf\it\mathcal{V}}$ (1)
@@
Here, ${\bf\it\mathcal{V}}$ is the set of kinematically admissible velocity fields, ${\bf\it\sigma}$ is the Cauchy stress, $\mathbf{D}$ is the symmetric rate of deformation tensor, $\mathbf{t}$ is the surface traction, $V$ is the total volume of the body, and $\partial V$ is the surface of the domain.
@@ 2.1.2 Downscaling
where $\mathbf{x}^{\text{s}}$ is the subscale coordinate, $\mathbf{F}^{\text{M}}$ is the macroscopic deformation gradient at the integration point, and $\tilde{\mathbf{x}}^{\text{s}}$ are the higher-order terms that are typically described as fluctuations. To proceed, we make the following choices, which can be interpreted as orthogonality of the mean and fluctuating fields [31, 32].
@@ 2.1.4 Subscale
In this work, we use Delaunay fiber networks to model the RVE. The network construction procedure proceeds as follows: First, a generation box is seeded with points using a uniform random distribution. A Delaunay triangularization is then constructed from the initial seed points. To avoid boundary effects, the generation box is trimmed so that the new edge length is half of the one used for the generation box. New nodes are inserted at the intersection of the trimming box and any crossing fibers. The fiber segments that remain outside the trimming box are deleted. This process will yield fiber networks with an average connectivity of $z\approx 14$. This is well above the isostaticity limit for 3D truss networks. The constitutive model for each fiber is given by:
@@
Figure 2: Schematic of the input convex neural network we implement. The input vector consists of invariants of the right Cauchy-Green deformation tensor ($\mathbf{C}$), and the output is the predicted strain energy density ($\mathcal{W}$). The components of the PK2 stress ($\mathbf{\Pi}$) and tangent stiffness tensor ($\mathcal{\mathbb{C}}=\partialderivative{\mathbf{\Pi}}{\mathbf{E}}$) are computed by back-propagation through the ICNN neural network. Cauchy stress $\sigma$ and the updated-Lagrangian material stiffness are needed in the finite element implementation and are computed through the appropriate push forward operations.
@@ 2.2.1 Model Selection
The hyperelastic formulation entails finding a non-negative, non-decreasing, and convex scalar energy density function ($\mathcal{W}$) in terms of deformation-related quantities [39, 40]. The stress and stiffness tensor components are then found through appropriate first and second-directional derivatives of $\mathcal{W}$. The convexity condition implies the machine learning algorithm must be chosen or designed appropriately [23, 41]. Here, we employ the Input Convex Neural Network (ICNN) model, which guarantees that the overall mathematical form of the neural network will be convex for appropriate choice of input and output quantities. The ICNN network was originally pioneered in [23]. Various authors have used the ICNN as a basis for hyperelastic material modeling [41, 17, 40, 24, 25]. However, this choice is not unique. Taç et al. provides an excellent benchmark between Constitutive Artificial Neural Networks (CANN), ICNNs, and Neural Ordinary Differential Equations, finding that they all exhibit significant limitations while being used to fit a skin dataset [42]. They additionally show that their trained networks exhibit large variations in the stiffness, further emphasizing the need to incorporate stiffness into the training dataset. Here, we chose to use the ICNN as we found our downstream finite element models were more likely to converge with the ICNN based network rather than the CANN network described in [25]. A schematic of the neural network model we implemented is shown in Figure 2.
2.2.2 Neural Network Architecture
@@
The output of a fully connected feed-forward neural network is a convex function of its input if the $i$-th layer of the neural network is a component-wise convex function of $i-1$-th layer [29]. ICNN is a particular form of fully connected feed-forward neural network with restrictions on activation functions and weights to satisfy this criterion. The convexity of the model is ensured by strictly adhering to (i) convex activation functions and (ii) non-negative weights. From an implementation point of view, the first requirement is trivial and entails using convex activation functions such as softplus [43, 29]. The second requirement, i.e., non-negative weights, limits the scope of the neural network when modeling complex phenomena [23]. To counter, a number of pass-through connections without restrictions on weights or biases are introduced that connect the inputs to the hidden and output layer neurons. These pass-through connections amount to an affine mapping of input and thus do not affect the convexity of the model.
[…]
@@
This model, however, does not guarantee vanishing strain energy at zero strain (i.e., $\mathcal{W}|_{\mathbf{F}=\mathbf{I}}$) due to the nature of activation functions. Specifically, the softplus activation function we use produces an output equals $g(0)=\log(2)>0$ at zero input, which leads to $\mathcal{W}|_{\mathbf{F}=\mathbf{I}}>0$. The optimization algorithm can overcome this shortcoming by adjusting the biases in the pass-through layers. Figure 4 indicates that the requirement is approximately satisfied. An alternative approach can be introducing standalone bias terms to act as correction factors, for example, in [17].
Finally, we note that the choice of input is dependent on the underlying material model and available dataset. For the isotropic and elastic fibrous material we model, our input array choice is appropriate. For anisotropic materials, additional quantities describing the anisotropy are suggested, e.g., [44]. Similarly, some invariants such as $I_{3}$ can be ignored if certain assumptions, such as incompressibility, are applicable. A brief discussion of the impact of removing the $I_{2}$ and $I_{3}$ invariants is presented in Section 3.2.
@@
Figure 3: Mean squared error (loss) of ICNN during training. The average errors associated with strain energy density, stress, and stiffness tensors are shown in the first, second, and third columns, respectively. The top row shows the loss for the training dataset, and the bottom row represents the corresponding results for the test dataset. The legend $\mathcal{L}^{k}:=\sum_{i=0}^{i=k}H_{i}$ indicates the variant of the loss function used to train ICNN, with $k$ being the highest order of derivative included in the composite loss function. The input array consists of all isotropic invariants.
2.2.3 ICNN Implementation
@@
The ICNN we train contains five fully connected hidden layers, each having four neurons in the primary branch. The pass-through layers connect the input vector to these hidden layer neurons. Mathematically, the output of the neurons in the $i$-th hidden layer is $\mathbf{z}^{(i)}=g(f(\mathbf{A}^{(i)})\mathbf{z}^{(i-1)}+\mathbf{A}^{(i)}_{p}\mathbf{x}+\mathbf{b}^{(i)}_{p})$. Here, $\mathbf{x}$ is the input vector to the model, and $A^{(i)}$ is the weight matrix for the $i$-th layer of the primary branch. The final two terms ($A^{(i)}_{p}$ and $b^{(i)}_{p}$) represent the weight matrix and bias vector due to the pass-through connection to the $i$-th layer, respectively. The model is trained and evaluated over 91,590 samples, with an 80/20 split for the train and test dataset. We use Adams solver for parameter optimization without any weight decay [45]. The model parameters are initialized using Kaiming uniform distribution [46]. We use the mean squared error $\text{MSE}(y)=(y_{\text{pred}}-y_{\text{true}})^{2}$ as the template for the loss function.
@@
In the original and most subsequent implementations of ICNN, the loss function is based only on the output variable. For our purpose, this translates to hyperparameter optimization based on strain energy density. However, due to the large number of trainable parameters, the prediction accuracy for energy does not generally translate to its derivative quantities, such as stress and stiffness, which are necessary for FE simulations. To increase the prediction accuracy of the derivatives, we implement the so-called Sobolev training protocol such that the model is trained on the target output and its derivatives concurrently [47]. Here, we include both the first and second derivatives in the ICNN training procedure as follows-
  1. 1.
    For the $i$-th batch iteration within an epoch, we compute the following average losses. The stress and stiffness components are found through automatic differentiation of ICNN updated at ($i-1$)th.
    1. (a)
      Loss in energy, $H_{0}=\text{MSE}(\mathcal{W})$
    2. (b)
      Loss in stress, $H_{1}=\langle\text{MSE}(\mathbf{\sigma})\rangle$ (averaged over the components)
    3. (c)
      Loss in stiffness, $H_{2}=\langle\text{MSE}(\mathop{\mathbb{D}})\rangle$ (averaged over the components)
  2. 2.
    The total loss per batch is computed as $\mathcal{L}^{k}=\sum_{j=0}^{j=k}\gamma_{j}H_{j}$, with $k=2$ being the highest order of derivative considered. Here, $\gamma_{j}$ is a weighting factor for each term. We find $\gamma_{j}=1\;\forall j\in[0,k]$ to be adequate despite the relative difference in units and magnitudes of $H_{j}$.
  3. 3.
    The model parameters are updated based on the total loss, and the cycle is repeated.
  4. 4.
    The learning rate is adjusted automatically at the end of an epoch based on the mean loss. The learning rate decreases by a factor of ten if the model exhibits non-decreasing loss for five consecutive epochs.
We use a moderately large batch size (128), so the loss (especially in the gradients) approximated from a batch is representative of the population. Finally, we note that the construction of neural network determines the convexity and smoothness of the model. Sobolev training is an add-on that greatly improves the prediction accuracy of higher-order quantities.
Sobolev training requires accurate first and second derivative calculations. To compute them, we follow the following procedure starting with the assumption that the underlying material behavior is hyperelastic (has a strain energy density function) which leads to the definition of the second Piola-Kirchhoff as $\Pi=\partialderivative{\mathcal{W}{}}{\mathbf{E}}=2\partialderivative{\mathcal{W}{}}{\mathbf{C}}$, where $\mathbf{C}$ is the right Cauchy-Green deformation tensor. And, the material stiffness is given as $\mathcal{\mathbb{C}}=\partialderivative{\mathbf{\Pi}}{\mathbf{E}}=2\partialderivative{\mathbf{\Pi}}{\mathbf{C}}$. The standard push forward operations are used to transform the stress and stiffness to the reported measures that correspond to those used by MuMFiM’s updated-Lagrangian finite element solution procedure. The push forward for the stiffness tensor is given in equation (6).
These derivatives of the strain energy density, $\mathcal{W}$, are computed through backpropagation and make use of the PyTorch auto differentiation capabilities. Although we use the invariants of the right Cauchy-Green deformation to restrict ourselves to an isotropic material model, they are computed using PyTorch vectors, which allows us to perform backpropagation to the right Cauchy-Green deformation tensor. Alternatively, we could have computed the derivatives with respect to the invariants and used the standard relations to obtain the stiffness and stress.
The performance of the backpropagation, especially for second derivatives, was significantly impacted by the implementation details. Details of these methods and ICNN implementations are provided in our GitHub repository  [^1] https://github.com/LACES-LAB/ML-biotissues/blob/main/imports/fibernn/derivatives.py and SI.
@@ 2.2.4 Training Data Generation
If a constant number of samples is used within each ellipsoidal shell that is represented by a given base stretch $\mathbf{U}^{*}_{i}$, then each data point will cover a varying volume in phase space. To avoid this problem, the number of rotation samples constructed for each set of base stretches is scaled by the surface area of the ellipsoid, with $\lambda_{1},\lambda_{2},\lambda_{3}$ forming the ellipsoid axes. This will ensure that, in a statistical sense, each sampled point will cover a constant volume in phase space, which is critical to avoid oversampling and associated overfitting. These sampled deformations are used as inputs to our GPU accelerated network solver in MuMFiM to generate the neural network training and test dataset. Some key characteristics of the generated dataset are provided in SI.
@@ 2.2.5 Model Availability
Figure 4: True (MuMFiM) vs ICNN-predicted strain energy density for various deformation gradients ($F$) in test dataset using ICNN trained with $\mathcal{L}^{2}$ loss function. The dashed red line indicates the ideal fit, and scatters indicate predictions.
@@ 3 Results and Discussions
In this section, we first present the model’s predictive capability and the improvements brought forth by the Sobolev training protocol. We then present a case study involving mechanical deformation of the FCL, simulated using the trained model. Finally, we compare the output to the results of MuMFiM based large-scale numerical simulation to comment on the efficacy of the current approach.
@@ 3.1 Sobolev Training of ICNN
We first discuss the overall predictive accuracy of the input convex neural network (ICNN) in terms of the average loss of desired quantities. This energy, stress, and stiffness result is shown in the first, second, and third columns of Figure 3. We compare three variants of the model emanating from the choice of the composite loss function (i.e., $k=0,1,2$ in $\mathcal{L}^{k}=\sum_{j=0}^{j=k}\gamma_{j}H_{j}$), indicating the highest derivatives included in the loss function.
The first column of Figure 3 shows the loss in strain energy density for different choices of the overall loss function. Expectedly, the choice of loss function has a negligible effect on the predictive accuracy of strain energy density. The derivative quantities (such as stress), however, benefit from including $H_{1}$ in the loss function (curves corresponding to $\mathcal{L}^{1}$). This result is shown in the second column of Figure 3. The additional information passed to the optimizer also improved the predictive accuracy of stiffness, as shown in the final column of the figure. Naturally, including the loss in stiffness in the overall loss function (curves corresponding to $\mathcal{L}^{2}$) improved the predictive accuracy of the stiffness for both the training and test datasets.
Comparing the loss function variants, it is evident that including higher-order derivatives of the output function improved the accuracy of the derivative quantities at any given epoch. The substantial improvement in stiffness prediction accuracy is especially beneficial for the downstream finite element algorithm, which would be a typical use case of the ICNN-based model developed here.
In fact, we found that the model variant with $\mathcal{L}^{2}$ led to a more robust finite element solution that was able to converge with a wider range solver parameters. And, the accuracy of the stresses was lower in the $\mathcal{L}^{0}$ and $\mathcal{L}^{1}$ cases. This impact is particularly noticable outside of range of the training data which can be seen in the context of a single RVE in Figure 11. Consequently, subsequent discussions primarily focus on the $\mathcal{L}^{2}$ variant only.